# An Allele-Centric Pan-Graph-Matrix Representation for Scalable Pangenome Analysis

### v1

@@ 1 An Allele-Centric Pan-Graph-Matrix Representation
The H1 pan-graph-matrix is not simply another instance of a binary matrix applied to genomic data. Its novelty lies in a redefinition of the unit of representation and in the adaptive manner in which population-scale genomic relationships are encoded. The contribution of H1 emerges at the intersection of pangenome modeling, compression-aware data representation, and population-level genomic analysis, where existing representations exhibit structural limitations [Paten2017GenomeGraphs, Eizenga2020PangenomeGraphs]. While recent work has substantially expanded catalogs of structural variation and enabled graph-based human pangenome references [Chaisson2019SV, Liao2023DraftHumanPangenome], scalable representations that make population-level allele incidence explicit remain an open challenge.
@@
Most existing genomic data representations implicitly fix both the semantic unit and the encoding scheme. Variant-centric formats such as VCF or BCF organize data by genomic position and record genotypes per sample using a rigid layout [Danecek2011VCF]. While effective for sample-centric queries, these formats are poorly aligned with population-level questions that require enumerating carriers of specific alleles. In contrast, H1 stores alleles as rows and provides direct access to the set of haplotypes carrying a given allele. As a consequence, compression efficiency in H1 follows population incidence patterns rather than file format constraints. Dense representations can also be implemented using compressed bitmap schemes [Chambi2016Roaring], which reduce constant factors but do not alter the underlying dense–sparse trade-off that governs the choice of representation.
H1 also differs fundamentally from haplotype-centric compression methods such as those based on the positional Burrows–Wheeler transform. These approaches are optimized for compressing entire haplotypes and excel in tasks such as phasing and imputation [Durbin2014PBWT, Li2016BGT], but they make allele-centric queries difficult to express efficiently. This limitation is particularly relevant for structural variants, whose population-scale diversity and functional impact have been highlighted by large multi-platform studies [Chaisson2019SV]. H1 deliberately avoids compressing haplotypes and instead compresses the incidence relation between alleles and haplotypes, shifting the representation from a path-centric to a relation-centric view. This distinction explains why H1 remains effective across both single-nucleotide and structural variants and complements existing haplotype compressors.
Graph-based pangenome representations encode genomic alternatives through nodes, edges, and paths, making structural rearrangements explicit and supporting alignment and traversal operations [Garrison2018VG, Li2020MinigraphrGFA]. However, population structure is implicit in these graphs, and compression is primarily driven by sequence redundancy rather than by carrier sparsity. The pan-graph-matrix can be understood as the incidence algebra of the pangenome graph: each alternative path corresponds to one or more rows of the matrix, and each haplotype path corresponds to a column. Both representations are equivalent in information content with respect to allele–haplotype incidence and ordered haplotype traversals, but emphasize different analytical dimensions.
@@ 2 Matrix–Graph Correspondence
All experiments were conducted on phased SNV/INDEL and structural variant callsets from the 1000 Genomes Project 30× high-coverage sequencing data [ByrskaBishop2022HighCoverage]. Quantitative compression results for the same genomic region are reported in Table 2. For a 2 Mb window on chromosome 1 comprising 200 diploid individuals (400 haplotypes), single-nucleotide and short insertion–deletion variants account for 24,921 sites, while only 45 sites correspond to true structural variants after filtering by SVTYPE. Despite their small number, structural variants exhibit an extreme sparsity regime: more than half are carried by two or fewer haplotypes, and nearly 87% have allele frequency below 10%.
@@ 2.2 Operations enabled by H1 and H2
Beyond storage, the H1 and H2 representations support a small set of primitive operations that are central to population-scale pangenome analysis. The cost of these operations depends primarily on the carrier count $k$ of an allele and the total number of haplotypes $H$. Query-oriented genotype stores [Layer2016GQT] optimize specific access patterns over large cohorts but retain sample- or haplotype-centric layouts, whereas H1 and H2 expose allele incidence and path structure directly. Further, these operations illustrate how H1 emphasizes population incidence and set-based queries, while H2 restores ordering and topological context. Together, they support a range of analytical tasks without conflating population-level incidence with path structure.
@@ 4 Limitations and Scope
This preprint focuses on representational structure and compression behavior. Formal derivations, broader scaling experiments, and downstream application benchmarks are left to future work.

### v2

@@ 1 An Allele-Centric Pan-Graph-Matrix Representation
The H1 pan-graph-matrix is not simply another instance of a binary matrix applied to genomic data. Its novelty lies in a redefinition of the unit of representation and in the adaptive manner in which population-scale genomic relationships are encoded. The contribution of H1 emerges at the intersection of pangenome modeling, compression-aware data representation, and population-level genomic analysis, where existing representations exhibit structural limitations [12, 6]. While recent work has substantially expanded catalogs of structural variation and enabled graph-based human pangenome references [2, 11], scalable representations that make population-level allele incidence explicit remain an open challenge.
@@
Most existing genomic data representations implicitly fix both the semantic unit and the encoding scheme. Variant-centric formats such as VCF or BCF organize data by genomic position and record genotypes per sample using a rigid layout [4]. While effective for sample-centric queries, these formats are poorly aligned with population-level questions that require enumerating carriers of specific alleles. In contrast, H1 stores alleles as rows and provides direct access to the set of haplotypes carrying a given allele. As a consequence, compression efficiency in H1 follows population incidence patterns rather than file format constraints. Dense representations can also be implemented using compressed bitmap schemes [3], which reduce constant factors but do not alter the underlying dense–sparse trade-off that governs the choice of representation.
H1 also differs fundamentally from haplotype-centric compression methods such as those based on the positional Burrows–Wheeler transform. These approaches are optimized for compressing entire haplotypes and excel in tasks such as phasing and imputation [5, 9], but they make allele-centric queries difficult to express efficiently. This limitation is particularly relevant for structural variants, whose population-scale diversity and functional impact have been highlighted by large multi-platform studies [2]. H1 deliberately avoids compressing haplotypes and instead compresses the incidence relation between alleles and haplotypes, shifting the representation from a path-centric to a relation-centric view. This distinction explains why H1 remains effective across both single-nucleotide and structural variants and complements existing haplotype compressors.
Graph-based pangenome representations encode genomic alternatives through nodes, edges, and paths, making structural rearrangements explicit and supporting alignment and traversal operations [7, 10]. However, population structure is implicit in these graphs, and compression is primarily driven by sequence redundancy rather than by carrier sparsity. The pan-graph-matrix can be understood as the incidence algebra of the pangenome graph: each alternative path corresponds to one or more rows of the matrix, and each haplotype path corresponds to a column. Both representations are equivalent in information content with respect to allele–haplotype incidence and ordered haplotype traversals, but emphasize different analytical dimensions.
@@ 2 Matrix–Graph Correspondence
All experiments were conducted on phased SNV/INDEL and structural variant callsets from the 1000 Genomes Project 30× high-coverage sequencing data [1]. Quantitative compression results for the same genomic region are reported in Table 2. For a 2 Mb window on chromosome 1 comprising 200 diploid individuals (400 haplotypes), single-nucleotide and short insertion–deletion variants account for 24,921 sites, while only 45 sites correspond to true structural variants after filtering by SVTYPE. Despite their small number, structural variants exhibit an extreme sparsity regime: more than half are carried by two or fewer haplotypes, and nearly 87% have allele frequency below 10%.
@@ 2.2 Operations enabled by H1 and H2
Beyond storage, the H1 and H2 representations support a small set of primitive operations that are central to population-scale pangenome analysis. The cost of these operations depends primarily on the carrier count $k$ of an allele and the total number of haplotypes $H$. Query-oriented genotype stores [8] optimize specific access patterns over large cohorts but retain sample- or haplotype-centric layouts, whereas H1 and H2 expose allele incidence and path structure directly. Further, these operations illustrate how H1 emphasizes population incidence and set-based queries, while H2 restores ordering and topological context. Together, they support a range of analytical tasks without conflating population-level incidence with path structure.
@@ 3 Graph-based visualizations of the H1 pangenome
4 Structure-Aware Compression of Graph Paths
Haplotype paths in pangenome graphs are not arbitrary symbol sequences: they are constrained walks over a directed graph with strong local dependencies. Exploiting these dependencies is essential to approach the information-theoretic lower bound on compressibility. The implemented compressor explicitly models each path as a realization of a first-order Markov process over graph nodes, capturing local transition regularities without requiring global enumeration of all possible paths.
4.1 Entropy-Optimal Coding with Minimal Overhead
By employing context-dependent Huffman coding, the compressor achieves expected code lengths within one bit of the Shannon entropy for each contextual distribution. Unlike naïve per-edge encodings, this approach avoids systematic over-penalization of high-probability transitions, which are prevalent in real haplotype data. At the same time, Huffman coding avoids the numerical instability and implementation complexity associated with arithmetic coding or ANS, while still providing strong entropy guarantees. This makes the method particularly well suited to sparse, skewed distributions typical of pangenome transitions.
4.2 Sparsity Preservation and Scalability
A key design choice is to restrict each context alphabet to its empirically observed support. This has two important consequences:
  • the cost of building and storing codebooks scales with observed transitions rather than with the total number of graph nodes;
  • memory usage remains linear in the number of distinct transitions present in the dataset.
This sparsity-aware construction is critical for large H2 instances, where full-alphabet models would be prohibitively expensive and largely uninformative.
4.3 Self-Delimiting and Robust Encoding
The explicit introduction of a STOP symbol ensures that encoded paths are self-delimiting and can be decoded without auxiliary length information. This property is essential for streaming, random access, and partial decoding scenarios. Moreover, the optional escape mechanism provides a controlled extension path for handling unseen transitions, preserving decodability even when the empirical model is imperfect or incrementally updated.
4.4 Practical Advantages over Alternative Methods
Compared to arithmetic coding or ANS-based approaches, the adopted compressor offers:
  • deterministic decoding without floating-point arithmetic;
  • bounded per-symbol redundancy, independent of total path length;
  • simpler correctness guarantees and easier debugging;
  • compatibility with partial decoding and diagnostic analysis.
Compared to naïve Markov or edge-wise encoders, it achieves substantially lower bit rates by adapting code lengths to empirical transition probabilities.
5 Implementation
We implement an entropic compressor for H2 haplotype paths based on a first-order Markov model with context-dependent Huffman coding. The design follows the Conditional Heavy Components Decomposition (CHCD) philosophy for general graphs, but specializes it to a sparse, empirical Markov setting suitable for pangenome haplotype paths [13]. Each haplotype path is encoded as a sequence of node indices, terminated by an explicit end-of-path symbol. Compression proceeds by modeling the conditional distribution of the next node given the current node (or a distinguished start context), and entropy-coding each transition with a locally optimal prefix code.
5.1 Graph and Path Normalization
H2 paths may be represented either as sequences of edge identifiers or directly as node labels. To ensure uniform processing, we first normalize all representations into sequences of node indices.
Given an edge map $E$ describing directed edges $(u,v)$, we construct:
  • a deterministic mapping from node labels to contiguous indices;
  • an adjacency list representation with duplicate edges removed;
  • a conversion routine that expands edge-id paths into node-label paths and then into node-index paths.
This normalization ensures that all downstream probabilistic modeling operates on a compact integer alphabet, independent of the original H2 encoding.
5.2 Markov Model with Explicit STOP Symbol
Let $\mathcal{P}$ be the set of observed haplotype paths. Each path $p=(x_{1},x_{2},\dots,x_{k})$ is modeled as a realization of a first-order Markov chain with:
  • a distinguished start context START,
  • an explicit terminal symbol STOP.
Empirical transition counts are collected as:
$\texttt{START}\rightarrow x_{1},\quad x_{i}\rightarrow x_{i+1},\quad x_{k}\rightarrow\texttt{STOP}.$
This construction ensures that:
  1. 1.
    all paths are self-delimiting;
  2. 2.
    decoding is unambiguous without external length metadata;
  3. 3.
    entropy is computed over complete paths rather than prefixes.
5.3 Sparse Context Alphabets and Additive Smoothing
For each context $c$ (either START or a node index), we observe a sparse set of successor symbols. Let $C_{c}(s)$ denote the empirical count of symbol $s$ following context $c$.
We apply additive (Laplace) smoothing with parameter $\alpha$:
$w_{c}(s)=C_{c}(s)+\alpha,$
restricted to the observed support of each context, with optional inclusion of:
  • STOP, forced into every context alphabet;
  • ESC, an escape symbol for unseen transitions (disabled by default).
This yields well-defined conditional distributions while preserving sparsity and avoiding the combinatorial explosion of a full alphabet.
5.4 Context-Dependent Huffman Coding
For each context $c$, a Huffman code is constructed from the smoothed weights $\{w_{c}(s)\}$. The resulting code minimizes the expected code length
$L_{c}=\sum_{s}p_{c}(s)\,\ell_{c}(s),$
subject to the prefix-free constraint, and satisfies the classical bound:
$H_{c}\leq L_{c}<H_{c}+1,$
where $H_{c}$ is the entropy of the smoothed conditional distribution.
To avoid rebuilding identical codebooks, each context is summarized by a signature consisting of:
  • the support of its counter,
  • the smoothing and configuration flags.
Contexts sharing the same signature reuse a cached Huffman tree, yielding substantial memory and time savings in large H2 instances.
5.5 Encoding and Decoding
Encoding proceeds sequentially:
  1. 1.
    emit the codeword for $x_{1}$ under context START;
  2. 2.
    for each transition $x_{i}\rightarrow x_{i+1}$, emit the context-specific codeword;
  3. 3.
    emit the STOP symbol from the final context.
Decoding mirrors this process using per-context decoding tries. Because all codes are prefix-free and STOP is explicit, decoding terminates deterministically without side information.
5.6 Entropy Tightness Diagnostics
To assess compression optimality, we compute for each context $c$:
  • the empirical entropy $H_{c}$ of the smoothed distribution;
  • the expected Huffman length $L_{c}$;
  • the gap $L_{c}-H_{c}$.
Aggregating over contexts weighted by empirical usage yields a global estimate:
$H_{\text{global}}\leq L_{\text{global}}<H_{\text{global}}+1,$
as predicted by Shannon–Huffman theory.
We additionally report the realized bits per symbol over the full dataset, verifying that observed compression closely tracks theoretical entropy and that no pathological contexts violate the expected bounds.
5.7 Correctness and Practical Guarantees
The implementation enforces:
  • exact round-trip correctness (encode followed by decode);
  • strict prefix-freeness in every context;
  • bounded redundancy (<1<1 bit per symbol per context).
This establishes the compressor as a faithful, efficient instantiation of CHCD-style entropic path compression specialized to H2 haplotype graphs.
5.8 Computational Performance and Compression Efficiency
We evaluate the performance of the proposed H2 entropic compressor along three dimensions: computational complexity, realized compression rate, and entropy tightness.
Computational Complexity.
Let $N$ be the number of haplotype paths and $L$ the average path length. Model fitting consists of a single pass over all paths to collect transition counts, followed by the construction of one Huffman code per distinct context. Since each context alphabet is sparse, with size equal to the number of empirically observed successors, the total fitting cost is
$\mathcal{O}\!\left(\sum_{p\in\mathcal{P}}|p|\;+\;\sum_{c}|\Sigma_{c}|\log|\Sigma_{c}|\right),$
where $\Sigma_{c}$ is the successor set of context $c$. In practice, $|\Sigma_{c}|$ is small, and identical contexts frequently share the same support; caching of Huffman codebooks further reduces overhead.
Encoding and decoding of a single path are linear in path length:
$\mathcal{O}(|p|),$
since each symbol emission or decoding step is a constant-time traversal of a context-specific Huffman table or trie.
Realized Compression Rate.
We measure the effective compression rate as the realized number of bits per emitted symbol:
$\bar{L}_{\text{real}}=\frac{\text{total encoded bits}}{\text{total emitted symbols (incl.\ STOP)}}.$
Empirically, this value closely matches the weighted expected code length derived from the fitted model, confirming that the compressor behaves predictably on real H2 data and that overhead from context switching or termination symbols is negligible.
Entropy Tightness.
For each context $c$, we compare the Shannon entropy $H_{c}$ of the smoothed conditional distribution with the expected Huffman length $L_{c}$. By construction, the Huffman code satisfies
$H_{c}\leq L_{c}<H_{c}+1.$
Aggregating contexts weighted by empirical usage yields global measures $H_{\mathrm{global}}$ and $L_{\mathrm{global}}$, for which we observe
$L_{\mathrm{global}}-H_{\mathrm{global}}\ll 1\;\text{bit/symbol}.$
No context exhibits a redundancy exceeding the theoretical Huffman bound, and high-frequency contexts dominate the global average, resulting in near-entropy-optimal compression.
5.9 Experimental Evaluation
We evaluate the proposed sparse Markov–Huffman compressor on a real pangenome dataset to assess its empirical compression behavior and its proximity to information-theoretic limits under realistic conditions.
Dataset.
[…]
Experimental Setup.
Haplotype paths are encoded using the proposed first-order Markov compressor with context-dependent Huffman coding. Transition probabilities are estimated empirically from the full dataset using additive smoothing ($\alpha=1$), and an explicit STOP symbol is used to ensure self-delimiting encodings. Compression performance is measured in terms of entropy, expected code length, and realized bits per emitted symbol. To avoid unstable estimates, contexts supported by fewer than ten observations are excluded from aggregate statistics.
Compression Ratio.
To quantify space reduction, we compare the compressed bitstream against a conservative baseline in which each emitted symbol (node index or STOP) is stored as a fixed-width 32-bit integer. Let $\bar{L}_{\mathrm{real}}$ denote the realized number of bits per symbol. The resulting compression ratio is
$\text{Compression ratio}=\frac{32}{\bar{L}_{\mathrm{real}}}.$
In our experiment, $\bar{L}_{\mathrm{real}}=1.0218$, yielding a compression ratio of approximately $31.3\times$, consistent with near-entropy-limited behavior.
Beyond per-symbol rates, we report the effective on-disk compression ratio of the serialized output. The uncompressed data occupy 81,602,97481{,}602{,}974 bytes ($\approx 77.8$ MiB), while the compressed file requires 5,567,7825{,}567{,}782 bytes ($\approx 5.31$ MiB), yielding a $14.66\times$ reduction. This ratio includes full model overhead (codebooks and indexing), and is therefore lower than the idealized bits-per-symbol estimate, but still reflects substantial savings under realistic storage conditions.
Table 3: Compression results for H2 haplotype paths using the sparse Markov–Huffman (CHCD-style) compressor.
Dataset / File Value
[…]
Total emitted symbols 20,319,698
Contexts (min. events $\geq 10$) 63,566
Global entropy $H_{\text{global}}$ (bits/symbol) 0.0762
Expected Huffman length $L_{\text{global}}$ (bits/symbol) 1.0253
Realized bits per symbol 1.0218
Redundancy LHL-H (bits/symbol) 0.9491
Compression ratio (vs. 32-bit symbols) $31.30\times$
Compression ratio on-disk (raw/cmp) $14.66\times$
Results.
Table 3 reports global compression statistics aggregated over all retained contexts. The realized compression rate closely tracks the expected Huffman code length and satisfies the theoretical bound $H\leq L<H+1$. When compared against a baseline representation using 32-bit symbols, the proposed method achieves a compression ratio of approximately $31\times$, indicating substantial space savings despite highly sparse and skewed transition distributions.
Discussion.
Together, these results indicate that the proposed compressor achieves: (i) encoding and decoding time linear in the length of each haplotype path, (ii) compression rates tightly concentrated around the empirical entropy, and (iii) robustness to sparsity and highly skewed transition distributions. The observed empirical behavior is consistent with CHCD-style guarantees for Markov path compression on general graphs, while remaining simple enough for practical deployment on large pangenome H2 instances [13].
@@ 6 Limitations and Scope
This preprint focuses on the representational structure and compression behavior of the proposed method. Formal theoretical derivations, large-scale benchmarking, and systematic evaluation on downstream applications are beyond the scope of the present work and are left for future investigation.
Appendix
Appendix A Pseudo-code: Sparse Markov–Huffman Compressor
Input: H2 object $h2$ with haplotype paths; edge map $E$ (optional); smoothing $\alpha\geq 0$; flags force_stop, enable_escape
Output: Model $\mathcal{M}$ containing per-context Huffman codes and decode tries
Normalize paths
$node\_to\_idx,idx\_to\_node,out\_adj\leftarrow\textsc{BuildGraphFromEdgeMap}(E)$
$paths\leftarrow\textsc{ExtractHapPaths}(h2)$
$P\leftarrow\textsc{ToNodeIndexPaths}(paths,E,node\_to\_idx)$
Count transitions (sparse)
Initialize dictionary of counters: $Counts[c]\leftarrow\emptyset$ for all contexts $c$
foreach path $p=(x_{1},\dots,x_{k})$ in $P$ do
 $Counts[\texttt{START}][x_{1}]\mathrel{+}=1$
 for $i\leftarrow 1$ to k1k-1 do
    $Counts[x_{i}][x_{i+1}]\mathrel{+}=1$
 $Counts[x_{k}][\texttt{STOP}]\mathrel{+}=1$
Build codebooks with caching
Initialize cache map: $Cache[sig]\to(Code,Trie)$
foreach context $c$ in CountsCounts do
 $sig\leftarrow\textsc{Signature}(Counts[c],\alpha,\texttt{force\_stop},\texttt{enable\_escape})$
 if $sig\in Cache$ then
    $(Code[c],Trie[c])\leftarrow Cache[sig]$
 else
    $Weights\leftarrow\{s:Counts[c][s]+\alpha\ \mid\ s\in\text{keys}(Counts[c])\}$
    if force_stop then
       $Weights[\texttt{STOP}]\mathrel{+}=\alpha$ (if missing, treat count as 0)
    if enable_escape then
       $Weights[\texttt{ESC}]\mathrel{+}=\alpha$ (if missing, treat count as 0)
    $Code[c]\leftarrow\textsc{BuildHuffmanCode}(Weights)$
    $Trie[c]\leftarrow\textsc{BuildDecodeTrie}(Code[c])$
    $Cache[sig]\leftarrow(Code[c],Trie[c])$
Return model $\mathcal{M}=(Code,Trie,Counts,\alpha)$
Algorithm 1 Fit sparse Markov–Huffman model on H2 paths
Input: Model $\mathcal{M}$; path $p=(x_{1},\dots,x_{k})$
Output: Bitstring $B$
$B\leftarrow\epsilon$
// empty bitstring
$c\leftarrow\texttt{START}$
// Emit first symbol from START
$B\leftarrow B\ \|\ Code[c][x_{1}]$
$c\leftarrow x_{1}$
// Emit transitions
for $i\leftarrow 2$ to $k$ do
 $B\leftarrow B\ \|\ Code[c][x_{i}]$
 $c\leftarrow x_{i}$
// Emit STOP
$B\leftarrow B\ \|\ Code[c][\texttt{STOP}]$
Return $B$
Algorithm 2 Encode a haplotype path
Input: Model $\mathcal{M}$; bitstring $B$
Output: Decoded path $p$
$p\leftarrow[]$
$i\leftarrow 1$
// bit index
$c\leftarrow\texttt{START}$
while true do
 $node\leftarrow Trie[c]$
 while nodenode has no terminal symbol do
    if $i>|B|$ then
       error (truncated stream)
    $b\leftarrow B[i]$; $i\leftarrow i+1$
    if $b\notin node$ then
       error (invalid prefix)
    $node\leftarrow node[b]$
 $s\leftarrow node[\mathdollar$]
 // decoded symbol
 if $s=\texttt{STOP}$ then
    break
 if $s=\texttt{ESC}$ then
    error (escape side-channel not implemented)
  append $s$ to $p$
 $c\leftarrow s$
Return $p$
Algorithm 3 Decode a bitstring into a haplotype path
Input: Model $\mathcal{M}$; minimum events threshold $m$
Output: Weighted global entropy $H_{\mathrm{global}}$ and expected length $L_{\mathrm{global}}$
Compute empirical usage per context: $Events[c]\leftarrow\sum_{s}Counts[c][s]$
$C^{\star}\leftarrow\{c:Events[c]\geq m\}$
$T\leftarrow\sum_{c\in C^{\star}}Events[c]$
$H_{\mathrm{global}}\leftarrow 0$; $L_{\mathrm{global}}\leftarrow 0$
foreach $c\in C^{\star}$ do
 $w\leftarrow Events[c]/T$
 $Syms\leftarrow\text{keys}(Code[c])$
 // sparse coded alphabet
 $Weights[s]\leftarrow Counts[c][s]+\alpha$ for $s\in Syms$
   Normalize: $p_{c}(s)\leftarrow Weights[s]/\sum_{u\in Syms}Weights[u]$
 $H_{c}\leftarrow-\sum_{s\in Syms}p_{c}(s)\log_{2}p_{c}(s)$
 $L_{c}\leftarrow\sum_{s\in Syms}p_{c}(s)\cdot|Code[c][s]|$
 $H_{\mathrm{global}}\mathrel{+}=w\cdot H_{c}$
 $L_{\mathrm{global}}\mathrel{+}=w\cdot L_{c}$
Return $(H_{\mathrm{global}},L_{\mathrm{global}})$
Algorithm 4 Entropy tightness diagnostics (global weighted)